# Few-Shot Domain Adaptation for Named-Entity Recognition via Joint Constrained k-Means and Subspace Selection

### v1

@@ Abstract
Ayoub Hammal1   Benno Uthayasooriyar2,3   Caio Corro4 1Universite Paris-Saclay, CNRS, LISN 2Data Analytics Solutions, SCOR  3LMBA, CNRS, Université de Brest 4INSA Rennes, IRISA, Inria, CNRS, Université de Rennes
@@ 3.2 Weak Supervision via Ratio Constraint
where the minimum is set to $\infty$- ∞ if the search space is empty.
@@ 4.2 Optimization Algorithm
We follow an alternative convex search procedure where variables are visited in order ${\bm{A}}\to{\bm{C}}\to{\bm{U}}$. Minimizing over ${\bm{A}}$ requires to take into account for the projection when computing matrix ${\bm{D}}$ in Equation (8), i.e. we set ${D}_{ij}=\|{\bm{U}}^{\top}{\bm{X}}_{i}-{\bm{U}}^{\top}{\bm{C}}_{j}\|^{2}$. Optimization over ${\bm{C}}$ is left unchanged (Appendix D).
@@ 5 Related Work
In the case of NER, Fritzler et al. (2019) adapted the prototypical network of Snell et al. (2017). Yang and Katiyar (2020) rely on nearest-neighbor classification along with a meta-transition matrix learned from the source task but with simpler IO transitions. Unfortunately, the later requires to tune a temperature hyper-parameter on the target domain, which is not realistically possible in the few-shot setting where there is no development set. Das et al. (2022) extended this approach by fine-tuning on both of the source dataset and target support using a contrastive loss function. Closer to our work, Hou et al. (2020) rely on a target task specific linear projection as proposed by Yoon et al. (2019), but they cannot benefit from extra unlabeled data.
Ratio constraints. Using supervision signal in the E step has been known as posterior regularization in the case of generative models (Ganchev et al., 2010). However, generic application of this framework rely on costly gradient descent to compute the solution of the E step, whereas we propose an polynomial analytical solution for our case. Previous work on ratio constraint for k-means reduced the E step to transportation problems (Ng, 2000; Bradley et al., 2000) but rely on generic algorithms, whereas we propose an efficient algorithm that benefit from the structure of our ratio constraint.
@@ 6 Experiments
When evaluating in the few-shot settings, we use the train set and development set belonging to each support as unlabeled data. For k-means, we use 10 iterations of the alternative convex search procedure. We have one cluster per entity type (I clusters), and we fix the number of clusters for the O tag to 10. As we cannot assume to know the true ratio of O tags, we evaluate our approach with both under- and over-estimations.
@@ 7 Conclusion
We propose a novel weakly-supervised algorithm for the few-shot NER. We evaluate our approach on different scenarios and achieve state-of-the-art results. Future work could consider applying our approach to learning from partial labels Jin and Ghahramani (2002) and to transductive learning.

### v2

@@ Abstract
Ayoub Hammal1   Benno Uthayasooriyar2,3   Caio Corro4 1Université Paris-Saclay, CNRS, LISN 2Data Analytics Solutions, SCOR  3LMBA, CNRS, Université de Brest 4INSA Rennes, IRISA, Inria, CNRS, Université de Rennes
@@ 3.2 Weak Supervision via Ratio Constraint
where the minimum is set to $\infty$ if the search space is empty.
@@ 4.2 Optimization Algorithm
We follow an alternate convex search procedure where variables are visited in order ${\bm{A}}\to{\bm{C}}\to{\bm{U}}$. Minimizing over ${\bm{A}}$ requires to take into account for the projection when computing matrix ${\bm{D}}$ in Equation (8), i.e. we set ${D}_{ij}=\|{\bm{U}}^{\top}{\bm{X}}_{i}-{\bm{U}}^{\top}{\bm{C}}_{j}\|^{2}$. Optimization over ${\bm{C}}$ is left unchanged, see App. D.
@@ 5 Related Work
In the case of NER, Fritzler et al. (2019) adapted the prototypical network of Snell et al. (2017). Yang and Katiyar (2020) rely on nearest-neighbor classification along with a meta-transition matrix learned from the source task but with simpler IO transitions. Unfortunately, the later requires to tune a temperature hyper-parameter on the target domain, which is not realistically possible in the few-shot setting where there is no development set. Das et al. (2022) extended this approach by fine-tuning on both of the source dataset and target support using a contrastive loss function. Chen et al. (2023) proposed to augment the input with output label information, similarly to the GliNER approach Zaratiana et al. (2024). Closer to our work, Hou et al. (2020) rely on a target task specific linear projection as proposed by Yoon et al. (2019), but they cannot benefit from extra unlabeled data.
Ratio constraints. Using a weak supervision signal in the E step has been known as posterior regularization in the case of generative models (Ganchev et al., 2010). However, generic application of this framework rely on costly gradient descent to compute the solution of the E step, whereas we propose a polynomial analytical solution for our hard k-means case and an efficient approach based on iterative Bregman projections for the soft case,. Previous work on ratio constraints for k-means reduced the E step to transportation problems (Ng, 2000; Bradley et al., 2000) but rely on generic algorithms, whereas we propose an efficient algorithm that benefit from the structure of our ratio constraint.
@@ 6 Experiments
When evaluating in the few-shot settings, we use the train set and development set belonging to each support as unlabeled data. For k-means, we use 10 iterations of the alternate convex search procedure. Note that this step takes approximately 6 minutes on CPU when using the full CoNLL dataset. We have one cluster per entity type (I clusters), and we fix the number of clusters for the O tag to 10. As we cannot assume to know the true ratio of O tags, we evaluate our approach with both under- and over-estimations.
@@ 7 Conclusion
We propose a novel weakly-supervised algorithm for the few-shot NER. We evaluate our approach on different scenarios and achieve state-of-the-art results. Future work could consider applying our approach to learning from partial labels Jin and Ghahramani (2002) and to transductive learning Colombo et al. (2023).